# Optimistic Query Routing in Clustering-based Approximate Maximum Inner Product Search

### v1

@@
Abstract
Clustering-based nearest neighbor search is a simple yet effective method in which data points are partitioned into geometric shards to form an index, and only a few shards are searched during query processing to find an approximate set of top-$k$ vectors. Even though the search efficacy is heavily influenced by the algorithm that identifies the set of shards to probe, it has received little attention in the literature. This work attempts to bridge that gap by studying the problem of routing in clustering-based maximum inner product search (MIPS). We begin by unpacking existing routing protocols and notice the surprising contribution of optimism. We then take a page from the sequential decision making literature and formalize that insight following the principle of “optimism in the face of uncertainty.” In particular, we present a new framework that incorporates the moments of the distribution of inner products within each shard to optimistically estimate the maximum inner product. We then present a simple instance of our algorithm that uses only the first two moments to reach the same accuracy as state-of-the-art routers such as Scann by probing up to $50\%$ fewer points on a suite of benchmark MIPS datasets. Our algorithm is also space-efficient: we design a sketch of the second moment whose size is independent of the number of points and in practice requires storing only $O(1)$ additional vectors per shard.
1 Introduction
@@
$\mathcal{S}=\operatorname*{arg\,max\quad}_{u\in\mathcal{X}}^{(k)}\langle q,u\rangle.$ (1)
Unsurprisingly, it is often too difficult to solve this problem exactly within a reasonable time budget, especially as $m$ or $d$ increases. As such, the problem is often relaxed to its approximate variant, where we tolerate error in the retrieved set to allow faster query processing. This approximate version of the problem is aptly named Approximate Nearest Neighbor (ANN) search, whose effectiveness is characterized by the fraction of true nearest neighbors recalled in the retrieved set: $\lvert\tilde{\mathcal{S}}\cap\mathcal{S}\rvert/k$, where $\tilde{\mathcal{S}}$ is the set returned by an ANN algorithm.
1.1 Clustering-based ANN search
Algorithms that solve the ANN search problem efficiently and effectively come in various flavors, from trees (Bentley, 1975; Dasgupta and Sinha, 2015), LSH (Indyk and Motwani, 1998), to graphs (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019) and more. For a thorough treatment of this subject, we refer the reader to (Bruch, 2024).
The method relevant to this work is the clustering-based approach, also known as Inverted File (IVF) (Jégou et al., 2011), which has proven effective in practice (Auvolat et al., 2015; Babenko and Lempitsky, 2012; Chierichetti et al., 2007; Bruch et al., 2024; Douze et al., 2024). In this paradigm, data points are partitioned into $C$ shards using a clustering function $\zeta:\operatorname*{\mathbb{R}}^{d}\rightarrow[C]$ on $\mathcal{X}$. A typical choice for $\zeta$ is the KMeans algorithm with $C=\mathcal{O}(\sqrt{m})$. This forms the index data structure.
Accompanying the index is a routing function $\tau:\mathbb{R}^{d}\rightarrow[C]^{\ell}$. It takes a query $q$ and returns $\ell$ shards that are more likely to contain its nearest neighbors. A commonly-used router is defined as follows:
$\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\langle q,\mu_{i}\rangle,$ (2)
@@ 1.1 Clustering-based ANN search
Processing a query $q$ involves two subroutines. The first stage, which we call “routing,” obtains a list of $\ell$ shards using $\tau(q)$, and the subsequent step, which we refer to as “search,” performs ANN search over the union of the selected shards. While a great deal of research has focused on the latter step (Jégou et al., 2011; Ge et al., 2014; Wu et al., 2017; Andre et al., 2021; Kalantidis and Avrithis, 2014; Johnson et al., 2021; Norouzi and Fleet, 2013), the former step has received relatively little attention. In this work, we turn squarely to the study of the first step: Routing queries to shards.
1.2 The importance of routing
@@
The historical focus on the search step makes a great deal of sense. After all, even though the routing step narrows down the search space, often dramatically so, selected shards may nonetheless contain a large number of points. It is thus imperative that the search stage be efficient and effective.
@@
We argue that the oft-overlooked routing step is important in its own right. The first and obvious reason is that, the more accurately [^1] We quantify the accuracy of a router as the ANN recall in a setup where the search stage performs a linear scan over the identified shards. As such, the only source of ANN error is the routing procedure, rather than inner product computation. a router chooses a subset of shards, the fewer data points the search stage must examine. For example, if shards are balanced in size, access to an oracle router means that the search stage need only examine $m/C$ points to find the top-$1$ point.
@@
The second and more important reason pertains to scale. As collections grow in size and dimensionality, it is often infeasible to keep the entire index in memory, in spite of advanced compression techniques such as Product Quantization (Jégou et al., 2011). Much of the index must therefore rest on secondary storage—in particular, cheap but high-latency storage such as disk or blob storage—and accessed only when necessary. That line of reasoning has led to the emergence of disk-based graph indexes (Jayaram Subramanya et al., 2019; Singh et al., 2021; Jaiswal et al., 2022) and the like.
Translating the same rationale to the clustering-based paradigm implies that shards rest outside of the main memory, and that when a router identifies a subset of shards, the search stage must fetch those shards from storage for further processing. A more accurate router thus lowers the volume of data that must be transferred between storage and memory. Furthermore, because query processing is additionally bottlenecked by storage bandwidth and memory capacity, fetching fewer shards per query is a requisite for achieving a higher throughput.
@@
Interestingly, depending on operational factors such as query load, storage I/O bandwidth, data transfer rate, and memory utilization, it would be acceptable for the routing stage to be more computationally expensive as long as it identifies shards more accurately. Whereas the in-memory clustering-based paradigm requires routing to be highly efficient, the new trade-off space offered by its storage-based realization opens the door to more nuanced research.
1.3 Existing routers and the surprising role of optimism
We have argued that routing accuracy is increasingly relevant. Surprisingly, with the exception of one recent work that explores supervised learning-to-rank for routing (Vecchiato et al., 2024), the few existing unsupervised routers take the naïve form of Equation (2). In effect, they determine the relative potential “reward” of every shard with a point estimate.
Take Equation (2) as the most prominent example. What we refer to as the Mean router summarizes each shard with its mean point (i.e., $\mu_{i}$ for shard $i$). This is not an unreasonable choice as the mean is the minimizer of the variance, and is, in fact, natural if KMeans is the clustering algorithm $\zeta$.
Another common router, which we call NormalizedMean, belongs to the same family, but where shard representatives are the $L_{2}$-normalized means, rather than the unnormalized mean vectors:
$\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\langle q,\frac{\mu_{i}}{\lVert\mu_{i}\rVert_{2}}\rangle.$ (3)
This formulation is inherited from the familiar Spherical KMeans (Dhillon and Modha, 2001), which is identical to the standard KMeans iterative algorithm but, at the end of every iteration, cluster centroids are projected onto the unit sphere. Because we can assume that $\lVert q\rVert_{2}=1$ without loss of generality, it is easy to see that Equation (3) routes by the angle between $q$ and the mean vectors.
Intuitively, NormalizedMean seems appropriate for ANN search over a sphere: When norms do not affect the outcome of ANN search, then all that matters in Equation (1) is the angle between $q$ and data points, rendering it reasonable to rank shards by the angle between their mean and $q$.
@@ 1.3 Existing routers and the surprising role of optimism
Figure 1: (a) Top-$1$ recall versus the number of points probed on the Text2Image dataset where data points have varying norms; (b) Distribution of inner products between a shard and a sample query on GloVe. Overlaid are the scores computed by Mean and NormalizedMean routers.
@@
Intriguingly, in many circumstances that deviate from that situation, NormalizedMean tends to perform more accurately than the Mean router, as evidenced in Figure LABEL:sub@figure:motivation:mean_vs_normalized_mean. Because unpacking this phenomenon motivates our proposal, we take a brief moment to take a closer look.
@@
Consider a query point $q$ and a single shard $\mathcal{P}$ with mean $\mu$. We visualize in Figure LABEL:sub@figure:motivation:score_distribution the distribution of inner products between $q$ and every point in $\mathcal{P}$. Overlaid with that distribution is the inner product between $q$ and the mean of points in $\mathcal{P}$, as well as their normalized mean. We observe that $\langle q,\mu/\lVert\mu\rVert_{2}\rangle$ lands in the right tail of the distribution.
@@
That is not surprising, at least for the case where $\lVert u\rVert_{2}=1$ for all $u\in\mathcal{P}$. Clearly $\lVert\mu\rVert_{2}\leq 1$, so that $\langle q,\mu/\lVert\mu\rVert_{2}\rangle=\langle q,\mu\rangle/\lVert\mu\rVert_{2}\geq\langle q,\mu\rangle$; the NormalizedMean router amplifies the Mean router by a factor of $1/\lVert\mu\rVert_{2}$. What is interesting, however, is that the magnitude of this “boost” correlates with the variance of the data points within $\mathcal{P}$: The more concentrated $\mathcal{P}$ is around the direction of its mean, the closer the mean is to the surface of the sphere, so that NormalizedMean applies a smaller amplification to $\langle q,\mu\rangle$. The opposite is true when points are spread out. As a result, shards with a higher variance receive a larger lift by the NormalizedMean router!
@@
This has its downsides. First, it is hard to explain the behavior on point sets with varying norms. Second, as we observe in Figure LABEL:sub@figure:motivation:score_distribution, NormalizedMean aggressively overestimates. Nonetheless, the insight that a router’s score for a shard can be influenced by the shard’s variance is worth exploring.
@@
We can summarize our observation as follows. The estimate given by the NormalizedMean router paints an optimistic picture of what the maximum inner product between $q$ and points in $\mathcal{P}$ could be. Contrast that with the Mean router, which is naturally a conservative estimate.
@@
This work investigates the ramifications of that insight. In particular, the research question we wish to study is whether a more principled approach to designing optimistic routers can lead to more accurate routing decisions on vector sets with variable norms. The nature of this question is not unlike those asked in the online learning literature (Lattimore and Szepesvári, 2020), so it is not surprising that our answer to this question draws from the “principle of optimism in the face of uncertainty.”
1.4 Contributions and outline
@@
2 Routing by the Optimism Principle: our proposal
@@
2.1 Formalizing the notion of optimism
@@
Equipped with $\theta_{i}$’s, we route $q$ by sorting all shards by their estimated thresholds in descending order and subsequently selecting the top shards in the resulting ranked list. Our router, dubbed Optimist, is defined as follow:
$\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\theta_{i}.$ (4)
2.2 Understanding the routing behavior
@@
Suppose for a moment that we have the solution to Problem 1, and let us expand on the expected behavior of the Optimist router. It is easy to see that, when $\delta\rightarrow 0$, the optimal solution approaches $\theta_{i}=\langle q,\mu_{i}\rangle$. As such, if we wish to obtain the most conservative estimate of the maximum inner product between $q$ and points in $\mathcal{P}_{i}$, the routing procedure collapses to the Mean router.
@@
As $\delta\rightarrow 1$, $\theta_{i}$ becomes larger, rendering Equation (4) an enthusiastically optimistic router. At the extreme, the optimal solution is the maximum inner product itself.
@@
Clearly then, $\delta$ controls the amount of optimism one bestows onto the router. It is interesting to note that, when the data distribution is fully known, then $\delta=1$ is an appropriate choice: If we know the exact distribution, we can expect to be fully confident about the maximum inner product. On the other hand, when very little about the distribution is known (e.g., when all we know is the first moment of the distribution), then $\delta=0$ is a sensible choice. In effect, the value of $\delta$ is a statement on our knowledge of the underlying data distribution.

### v3

@@
Abstract.
Clustering-based nearest neighbor search is an effective method in which points are partitioned into geometric shards to form an index, with only a few shards searched during query processing to find a set of top-$k$ vectors. Even though the search efficacy is heavily influenced by the algorithm that identifies the shards to probe, it has received little attention in the literature. This work bridges that gap by studying routing in clustering-based maximum inner product search. We unpack existing routers and notice the surprising contribution of optimism. We then take a page from the sequential decision making literature and formalize that insight following the principle of “optimism in the face of uncertainty.” In particular, we present a framework that incorporates the moments of the distribution of inner products within each shard to estimate the maximum inner product. We then present an instance of our algorithm that uses only the first two moments to reach the same accuracy as state-of-the-art routers such as Scann by probing up to $50\%$ fewer points on benchmark datasets. Our algorithm is also space-efficient: we design a sketch of the second moment whose size is independent of the number of points and requires $\mathcal{O}(1)$ vectors per shard.
maximum inner product search, clustering-based nearest neighbor search, query routing, sketching, optimism principle
ccs: Information systems Retrieval models and ranking
1. Introduction
@@
(1) $\mathcal{S}=\operatorname*{arg\,max\quad}_{u\in\mathcal{X}}^{(k)}\langle q,u\rangle.$
Unsurprisingly, it is often too difficult to solve this problem exactly within a reasonable time budget, especially as $m$ or $d$ increases. As such, the problem is often relaxed to its approximate variant, where we tolerate error in the retrieved set to allow faster query processing. This approximate version of the problem is aptly named Approximate Nearest Neighbor (ANN) search, whose effectiveness is characterized by the fraction of true nearest neighbors recalled in the retrieved set: $\lvert\tilde{\mathcal{S}}\cap\mathcal{S}\rvert/k$, where $\tilde{\mathcal{S}}$ is the set returned by an ANN algorithm. We call this metric recall, or equivalently, accuracy.
1.1. Clustering-based ANN search
ANN algorithms come in various flavors, from trees (Bentley, 1975; Dasgupta and Sinha, 2015), LSH (Indyk and Motwani, 1998), to graphs (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019). Refer to (Bruch, 2024) for a thorough review of this subject. The method relevant to this work is the clustering-based approach, also known as Inverted File (IVF) (Jégou et al., 2011), which has proven effective in practice (Auvolat et al., 2015; Babenko and Lempitsky, 2012; Chierichetti et al., 2007; Bruch et al., 2024; Douze et al., 2024). In this paradigm, data points are partitioned into $C$ shards using a clustering function $\zeta:\operatorname*{\mathbb{R}}^{d}\rightarrow[C]$ on $\mathcal{X}$. A typical choice for $\zeta$ is the KMeans algorithm with $C=\mathcal{O}(\sqrt{m})$. This forms the index data structure.
Accompanying the index is a routing function $\tau:\mathbb{R}^{d}\rightarrow[C]^{\ell}$. It takes a query $q$ and returns $\ell$ shards that are more likely to contain its nearest neighbors. A common router is as follows:
(2) $\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\langle q,\mu_{i}\rangle,$
@@ 1.1. Clustering-based ANN search
Processing a query $q$ involves two independent subroutines. The first stage, which we call “routing,” obtains a list of $\ell$ shards using $\tau(q)$, and the subsequent step, “scoring,” computes inner product between the query and the points in the union of the selected shards. While a great deal of research has focused on scoring step (Jégou et al., 2011; Ge et al., 2014; Wu et al., 2017; Andre et al., 2021; Kalantidis and Avrithis, 2014; Johnson et al., 2021; Norouzi and Fleet, 2013), routing has received relatively little attention. This work turns squarely to the first step: Routing queries to shards.
Before we proceed further, we wish to emphasize the following point: In a clustering-based ANN system, we can select a routing protocol to choose shards, independently of the scoring algorithm that evaluates the chosen shards. For example, we may route queries to shards (i.e., decide which shards to probe) using Equation (2) or using our novel router to be introduced later. This choice has no algorithmic bearing on the scoring phase whatsoever. In other words, the scoring stage can perform a linear scan of the chosen shards, or it may compute inner products using Product Quantization (PQ) (Jégou et al., 2011), or it may further index and search each shard using a graph ANN algorithm (Jayaram Subramanya et al., 2019; Malkov and Yashunin, 2020). All these choices are independent of the type of router utilized in the earlier stage. This is indeed one of the advantages of a clustering-based ANN search system.
That independence simplifies the evaluation of different routing protocols: By independence and the additivity of latency, any efficiency savings achieved by a router translates directly into efficiency savings, of the same magnitude, for the end-to-end retrieval system (i.e., routing followed by scoring). For instance, if two routers lead to the same end-to-end recall, but one does so at a substantially lower latency, its latency savings will be reflected in its end-to-end query latency. As a result, we can fix the scoring stage across our experiments (to an exact linear scan, e.g.), and examine routers in isolation to compare their accuracy, latency, and other characteristics.
1.2. The importance of routing
@@
The historical focus on the scoring phase makes sense. Even though the routing step narrows down the search space, selected shards may nonetheless contain a large number of points. It is thus imperative that the scoring stage be efficient and effective.
@@
We argue that the oft-overlooked routing step is important in its own right. The first and obvious reason is that, the more accurately [^1] We quantify the accuracy of a router as the ANN recall in a setup where the scoring stage performs a linear scan over the identified shards. As such, the only source of ANN error is the routing procedure, rather than inner product computation. a router chooses a subset of shards, the fewer data points the scoring stage must examine. For example, if shards are balanced in size, access to an oracle router (i.e., one that identifies the shard with the true nearest neighbor) means that the scoring stage must only compute inner products for $m/C$ points to find the top-$1$ point.
@@
The second reason pertains to scale. As collections grow in size and dimensionality, it is often infeasible to keep the entire index in memory, in spite of compression techniques such as PQ. Much of the index must therefore rest on secondary storage—in particular, cheap but high-latency storage such as disk or blob storage—and accessed only when necessary. That line of reasoning has led to the emergence of disk-based graph indexes (Jayaram Subramanya et al., 2019; Singh et al., 2021; Jaiswal et al., 2022) and the like.
Translating the same rationale to the clustering-based paradigm implies that shards rest outside of the main memory, and that when a router identifies a subset of shards, the scoring stage must fetch those shards from storage for further processing. This paradigm has gained traction in real-world vector database systems. [^2] See, for example, https://turbopuffer.com/blog/turbopuffer and https://www.pinecone.io/blog/serverless-architecture/#Pinecone-serverless-architecture. This is the context in which we present our research: Storage-backed, clustering-based ANN search systems.
In this framework, a more accurate router lowers the volume of data that must be transferred between storage and memory. That has several consequences including a lower utilization of the storage bandwidth per query, as well as a lower consumption of in-memory cache space per query, thereby improving retrieval throughput.
1.3. An illustrative example
@@
Interestingly, depending on operational factors such as query load, storage I/O bandwidth, data transfer rate, and memory utilization, it would be acceptable for the routing stage to be more computationally expensive as long as it identifies shards more accurately. Let us consider an illustrative example to support this statement and, in fact, motivate our research.
Consider Natural Questions (Kwiatkowski et al., 2019) embedded with Ada-002333https://openai.com/index/new-and-improved-embedding-model/ denoted by Nq-Ada2. It contains $2.7$ million $1{,}536$-dimension points. When clustered with KMeans, each shard contains about $1{,}650$ points, and is $1$MB in size, if compressed with $4$-bit PQ codebooks—a common configuration.
Next, consider a commercial machine: AWS c5.xlarge (with $4$ vCPUs and $8$GB of memory). Using $4$ threads, transferring $4$MB of data from blob storage (hosted on Amazon’s own S3) to main memory has an average latency of $45$ milliseconds (ms). [^4] See https://github.com/dvassallo/s3-benchmark for an independent and comprehensive benchmark.
Given this setup, fetching $16$ Nq-Ada2 shards from blob storage and downloading them to the machine’s main memory takes $180$ms (P50 latency). On this same machine, routing using Equation (2) takes $0.76$ms. As such, so long as a more complex router can achieve the same accuracy as Equation (2) by fetching $16$ fewer shards, but by introducing a computational latency of no greater than roughly $179$ms, the overall efficiency of the query processor improves. As we will show later, our router on the same dataset with the same configuration takes only about $10.1$ms to route queries to shards, leading to a saving of $170.6$ms per query in the given example.
As this example illustrates, a storage-backed, clustering-based ANN search system offers a trade-off space between not just accuracy and speed, but also other efficiency factors such as I/O bandwidth, and memory and cache utilization. This opens the door to more nuanced research. Our work is a step in that direction.
1.4. Existing routers and the role of optimism
@@
Figure 1. (a) Top-$1$ recall vs. percentage of points probed on Text2Image where points have varying norms; (b) Distribution of inner products between a shard and a query on GloVe. Overlaid are scores computed by Mean and NormalizedMean.
As we argued, routing accuracy is increasingly relevant. Surprisingly, with the exception of one recent work that explores supervised learning-to-rank for routing (Vecchiato et al., 2024), the few existing unsupervised routers take the naïve form of Equation (2).
Take Equation (2) as the most prominent example. What we refer to as the Mean router summarizes each shard with its mean point (i.e., $\mu_{i}$ for shard $i$). This is not an unreasonable choice as the mean is the minimizer of the sum of squared deviations from the mean, and is, in fact, natural if KMeans is the clustering algorithm $\zeta$.
Another common router, which we call NormalizedMean, belongs to the same family, but its shard representatives are the $L_{2}$-normalized means, rather than the unnormalized mean vectors:
(3) $\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\langle q,\frac{\mu_{i}}{\lVert\mu_{i}\rVert_{2}}\rangle.$
This formulation is inherited from the familiar Spherical KMeans (Dhillon and Modha, 2001), which is identical to standard KMeans except that, at the end of every iteration, cluster centroids are normalized. Because we can assume that $\lVert q\rVert_{2}=1$ without loss of generality, it is easy to see that Equation (3) routes by the angle between $q$ and the mean vectors.
Intuitively, NormalizedMean seems appropriate for ANN search over a sphere: When norms are constant, all that matters in Equation (1) is the angle between $q$ and data points, rendering it reasonable to rank shards by the angle between their mean and $q$.
@@
Intriguingly, in many circumstances that deviate from that situation, NormalizedMean performs more accurately than the Mean router, as evidenced in Figure LABEL:sub@figure:motivation:mean_vs_normalized_mean. Let us unpack this phenomenon.
@@
Consider a query point $q$ and a single shard $\mathcal{P}$ with mean $\mu$. We visualize in Figure LABEL:sub@figure:motivation:score_distribution the distribution of inner products between $q$ and every point in $\mathcal{P}$. Overlaid is the inner product between $q$ and the mean of points in $\mathcal{P}$, as well as their normalized mean. We observe that $\langle q,\mu/\lVert\mu\rVert_{2}\rangle$ lands in the right tail of the distribution.
@@
That is not surprising for the case where $\lVert u\rVert_{2}=1$ for all $u\in\mathcal{P}$. Clearly $\lVert\mu\rVert_{2}\leq 1$, so that $\langle q,\mu/\lVert\mu\rVert_{2}\rangle=\langle q,\mu\rangle/\lVert\mu\rVert_{2}\geq\langle q,\mu\rangle$; the NormalizedMean router amplifies the Mean router by a factor of $1/\lVert\mu\rVert_{2}$. What is interesting, however, is that the magnitude of this “boost” correlates with the variance of the data within $\mathcal{P}$: The more concentrated $\mathcal{P}$ is around the direction of its mean, the closer the mean is to the surface of the sphere, making $\lVert\mu\rVert_{2}$ larger, so that NormalizedMean applies a smaller amplification to $\langle q,\mu\rangle$. The opposite is true when points are spread out: shards with a higher variance receive a larger lift by the NormalizedMean router!
@@
There are a few caveats. First, it is hard to explain the behavior on point sets with varying norms (i.e., where $\lVert u\rVert_{2}\neq 1$). Second, as we observe in Figure LABEL:sub@figure:motivation:score_distribution, NormalizedMean aggressively overestimates. Nonetheless, the insight that a router’s score for a shard can be influenced by the shard’s variance is worth exploring.
@@
Let us summarize our observation. The estimate given by NormalizedMean paints an optimistic picture of what the maximum inner product between $q$ and points in $\mathcal{P}$ could be. Contrast that with the Mean router, which is naturally a conservative estimate.
@@
We investigate the ramifications of that insight. In particular, the research question we wish to study is whether a more principled approach to designing optimistic routers can lead to more accurate routing decisions on vector sets with variable norms. The nature of this question is not unlike those asked in the online learning literature (Lattimore and Szepesvári, 2020), so it is not surprising that our answer to this question draws from the “principle of optimism in the face of uncertainty.”
1.5. Contributions and outline
@@
2. Routing by the Optimism Principle: our proposal
@@
2.1. Formalizing the notion of optimism
@@
Equipped with $\theta_{i}$’s, we route $q$ by sorting all shards by their estimated thresholds in descending order and selecting the top shards. Our router, dubbed Optimist, is defined as follow:
(4) $\tau(q)=\operatorname*{arg\,max\quad}^{(\ell)}_{i\in[C]}\theta_{i}.$
2.2. Understanding the routing behavior
@@
Suppose for a moment that we have the solution to Problem 1, and let us expand on the expected behavior of the Optimist router. It is easy to see that, when $\delta\rightarrow 0$, the optimal solution approaches $\theta_{i}=\max(\langle q,\mu_{i}\rangle,\langle q,M_{i}\rangle)$, where $M_{i}$ is the median of the $i$-th partition. That represents the most conservative estimate of the maximum inner product between $q$ and points in $\mathcal{P}_{i}$.
@@
As $\delta\rightarrow 1$, $\theta_{i}$ becomes larger, rendering Equation (4) an enthusiastically optimistic router. At the extreme, the optimal solution is the maximum inner product itself—the most optimistic we can get.
@@
Clearly then, $\delta$ controls the amount of optimism one bestows onto the router. It is interesting to note that, when the data distribution is fully known, then $\delta=1$ is an appropriate choice: If we know the exact distribution of inner products, we can expect to be fully confident about the maximum inner product. On the other hand, when very little about the distribution is known (e.g., when all we know is the first moment of the distribution), then $\delta=0$ is a sensible choice. In effect, the value of $\delta$ is a statement on our knowledge of the underlying data distribution.